How many solutions are there to the following system of equations?
-10x – 4y = 100
15x + 6y = 150

Answer by Alan3354(69443)

How many solutions are there to the following system of equations?
-10x – 4y = 100
15x + 6y = 150
----------------
-10x – 4y = 100 --> 5x + 2y = -50
15x + 6y = 150 --> 5x + 2y = 50
-------------
No solutions.